In the right triangle ABC, AC= 12, angle B= 30 degrees, Angle C= 90 degrees, and Angle A= 60 degrees.
What is the perimeter of the Triangle
: In the right triangle ABC, AC= 12, angle B= 30 degrees, Angle C= 90 degrees, and Angle A= 60 degrees.
What is the perimeter of the Triangle

Answer by Alan3354(1216)

In the right triangle ABC, AC= 12, angle B= 30 degrees, Angle C= 90 degrees, and Angle A= 60 degrees.
What is the perimeter of the Triangle
-------------------
The hypotenuse is AB.
The sin(30) = AC/AB
.5 = AC/AB
AB = AC*2 = 24
---------
sin(60) = BC/AC
sqrt(3)/2 = BC/24
BC = 12*sqrt(3)
Circumference = 12+24+12sqrt(3)
= 36+12sqrt(3)